# Beyond Flat Text: Dual Self-inherited Guidance for Visual Text Generation

### v1

@@ 4.4 Quantitative Analysis
As shown in Tab. 2, our method shows substantial improvements over the competitive methods in both English and Chinese across all levels. Notably, our method doesn’t require additional training and enhances the performance of GlyphControl and AnyText at the easy level. At the hard level, AnyText with our method yields results closest to the ground truth in the NED metric, with an improvement of approximately $10\%.$ over the baseline AnyText in both languages.
@@
In the absence of ground-truth images, we use the CLIP score [8] to assess the consistency between the prompt and the generated image. We compute the average cosine similarity between the prompt and the generated image, excluding the influence of the visual texts. As shown in Tab. 2, our method increases the accuracy of visual text generation without compromising the baseline performance. Additionally, our method improves the text-image similarity slightly, due to our design that minimizes conflicts between the image composition and the visual text. This leads to a more coherent layout for both the image and the text.
@@
Following [2], we conduct user study to comprehensively compare the generation results in Tab. 3. We create 27 input sets with varying levels of difficulty in visual text generation. For each input set, we provide the respondents with our input prompt and position mask and two image alternatives: our result and a baseline (in random order). We ask them to select which image has the best quality and has the most accurate visual text. The final score is calculated based on the times the respondent choose the approach averaged on the total prompts. We collect 126 user judgments from a diverse group composed of both experts in the field and individuals with no specific background and report the percentage of votes. As shown in Tab. 3, our method is the most favored among all the circumstances.
@@ 5 Ablation Study
Figure 5: Ablation Study Visualization. The first column shows accuracy improvement with SRB. The second column demonstrates that while SIB enhances accuracy, it compromises coherence. The third column highlights that combining both branches yields accurate, coherent visual texts. Finally, AdaIN combination in the fourth column further refines image quality.
In this section, we demonstrate the effectiveness of our components through an ablation study on the medium level. We use OCR accuracy (Sen.Acc and NED) and CLIP score as the main metrics, given their importance in evaluating visual text generation. All parameters are kept consistent with those outlined in Sec. 4.1. The analysis of each component will be presented in the following parts of this section.
@@
Table 4: Ablation study. The results validate the effectiveness of each component.
@@ 6 Conclusion
In this work, we push the boundary of visual text generation by tackling the challenge of creating complex visual texts. We present a dual-branch method STGen, which decomposes the visual text generation into 2 branches: Semantic and Structure. In Semantic Rectification Branch, our method leverages the rich semantic information composed in the latent when generating accurate visual texts in simple scenarios to guide the generation in complex scenarios. In the Structure Injection Branch, we polish the visual text structure using the structural information extracted from glyphs at the same time. For challenging scenarios, we break them down into simpler steps to manage effectively. This approach can be integrated seamlessly with the existing visual text generation model using a dedicated ControlNet for generation. To exhibit its superior performance, we conduct extensive experiments on our proposed benchmark, demonstrating its excellence. Our approach offers a straightforward yet effective way to enhance the capabilities of existing visual text generation models. We believe STGen sets the stage for real-world applications, enabling more accurate and sophisticated visual text generation.
Supplementary Material
Figure 6: Comparison of predicted $x_{0}$ under different inference steps in various masks. The first column is the AnyText intermediate results under a flat mask, AnyText can predict the structure of the texts at the very early stage of inference. The rest of the columns show intermediates in a slanted mask using AnyText and other operations using different methods. Our method can revise the originally wrong texts and harmonize the background.
7 Additional Background
Latent Diffusion Models.
The Latent Diffusion Models (LDM) denoise a noisy vector $z_{t}$ to $z_{0}$ that is mapped to an image $x_{0}$ through an autoencoder, based on a text prompt $y$. To sequentially remove the noise $\varepsilon$, a network $\varepsilon_{\theta}$ is trained to minimize the loss:
$\mathcal{L}=\mathbb{E}_{z_{0},y,\varepsilon\sim\mathcal{N}(0,1),t}||\varepsilon-\varepsilon_{\theta}(z_{t},t,c(y))||_{2}^{2},$ (8)
where the $c(y)$ is the conditioning embedding of the prompt $y$, $z_{t}$ is a noisy vector obtained by adding noise to $z_{0}$ according to the timestep $t$. During inference, provided a random noise vector $z_{T}$, the trained network iteratively removes the predicted noise to produce a latent $z_{0}$ for $T$ steps. Namely, we employ popular DDIM sampling [29] for inference:
$z_{t-1}=$ $\sqrt{\alpha_{t-1}}\underbrace{(\frac{z_{t}-\sqrt{1-\alpha_{t}\epsilon_{\theta}(z_{t})}}{\sqrt{\alpha_{t}}})}_{\text{predicted $z_{0}$}}$ (9)
$+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}\cdot\epsilon_{\theta}(z_{t})}.$
The DDIM Sampling in general combines the predicted $z_{0}$ and $z_{t}$ to obtain the $z_{t-1}$. We mainly analyze the predicted $x_{0}$ that is decoded from predicted $z_{0}$ in the rest of the section.
Adaptive Instance Normalization.
The Adaptive Instance Normalization (AdaIN) [10] is previously adopted in the task of style transfer. It substitutes mean and standard deviation of the activations from each CNN [13] filter of the original image with those of the style image as follows:
${\rm AdaIN}(x,y)=\sigma(y)(\frac{x-\mu(x)}{\sigma(x)})+\mu(y),$ (10)
where $x$ and $y$ stand for the activation of a CNN filter from the content image and style image, $\mu(\cdot)$ and $\sigma(\cdot)$ are channel-wise mean and standard deviation. Masui et al. [19] further prove that the AdaIN can be directly applied to the LDM without any additional training for style transfer.
8 Analysis of Predicted $x_{0}$
As shown in Fig. 6, AnyText can produce clear and intact visual texts in structure even in the first inference step. However, once tilted, the prediction becomes blurry and wrong as in the second column. The following columns show the effectiveness of our method. SRB can improve the structure. The SIB further improves the structure but produces visual texts that are not coherent with the background.
Figure 7: Results under different lambda. When $\lambda$ is -0.5, we obtain colorful and styled visual texts in the first column. When $\lambda$ is 0.5, we produce images with rich background. When $\lambda$ is 0, the method degrades to vanilla AnyText.
Language $\lambda$ $\rho$ Sen.ACC NED CLIP score
easy medium hard total easy medium hard total
English -0.5 0.5 71.33 35.51 6.14 44.88 87.23 62.98 22.79 64.11 0.3005
0.5 0.5 71.25 37.25 6.60 45.43 87.54 64.42 24.56 65.10 0.3027
0.5 0.25 71.37 37.11 5.32 45.12 87.10 63.80 21.19 63.85 0.3036
0.5 0.75 69.60 38.39 7.16 45.10 87.33 65.54 25.94 65.65 0.3006
0.5 1.25 61.48 35.16 6.33 40.18 85.38 63.73 25.44 64.11 0.3008
0.5 1.5 69.47 38.92 6.48 45.01 87.23 65.03 25.43 65.34 0.3009
0.5 1.75 69.19 38.17 6.56 44.70 87.18 65.31 25.40 65.39 0.3011
0.5 2.0 68.88 37.96 6.71 44.53 87.07 64.85 25.26 65.17 0.3009
Chinese -0.5 0.5 68.61 35.02 7.54 49.28 95.17 85.23 66.05 87.29 0.3067
0.5 0.5 69.22 35.58 8.55 49.96 95.37 85.85 68.77 88.08 0.3071
0.5 0.25 71.35 35.42 6.16 50.70 95.46 85.88 67.37 87.86 0.3076
0.5 0.75 66.12 37.79 10.66 49.05 94.54 85.77 70.38 87.89 0.3058
0.5 1.25 64.28 38.11 10.57 48.03 94.31 86.14 71.58 88.07 0.3060
0.5 1.5 63.85 37.05 9.83 47.40 94.24 85.52 70.26 87.64 0.3061
0.5 1.75 65.53 38.11 10.29 48.12 94.29 85.85 70.21 87.73 0.3064
0.5 2.0 65.02 35.91 10.11 47.88 94.44 85.37 69.78 87.63 0.3059
Table 5: Detailed sensitivity analysis for $[\lambda,\rho]$.
9 Additional Results
Here we provide additional results of our method. Our code will be released upon acceptance.
9.1 Additional Quantitative Results
Hyper-parameters Sensitivity
As shown in the Tab. 5, our method shows robustness in hyper-parameters change. The best performance is obtained when $\rho$ is 0.75 at the hard level. But the method achieves the best performance in the easy level of Chinese and English set when $\rho$ is 0.25. To ensure fairness, we choose the performance when $\rho$ is 0.5 which has a balanced performance for comparison.
9.2 Additional Qualitative Results
Choice of $\lambda$
We can observe in Fig. 7 that when $\lambda$ is in the range of $(0,0.5]$ we can achieve background with more diversity and in the range of $[-0.5,0)$ the diversity of texts is increased consistent with what we mentioned in Sec. 4.1.
Coherence with Background
As shown in Fig. 8, thanks to the rich semantic information contained in the latent with flat visual texts, we can generate visual texts that are perfectly blended in with the background. The text ‘Harvest’ in the bottom right is more harmonious with the background and is closest to the description “interwoven”.
Visual Text Diversity
As shown in the first row of Fig. 10, with proper control, our method can generate texts in various styles and colors. The second row shows that the generated visual texts remain to be accurate under 2 different image styles. The third row demonstrates the robustness of our method under different challenging masks. The fourth row shows our ability to generate various visual texts using the same prompt.
Additional Comparison with baseline.
We present more qualitative comparisons on simple slanted or curved situations in Fig. 11 and Fig. 12, and complex scenarios in Fig. 13. Our method can generate high-quality images with accurate visual texts in all the scenarios.
10 Details of User Study
An illustration of our user study is provided in Fig. 9. Since the caption in the English (LAION) set are often hard for users to understand (e.g. “carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel of carousel”). To produce meaningful prompts that can help the user understand, we employ LLMs to generate image prompts for user study. Here we provide a simplified template of our instructions and examples:
1.Task instruction
I am working on a text-to-image generation task. Each Prompt should include a scene description and contain one or two sets of words that need to be present in the image, enclosed in double quotes.
2. Examples
A raccoon stands in front of the blackboard with the words “Deep Learning” written on it …… 3.Trigger CoT reasoning ability of LLMs
Reasoning: Let’s think step by step……
11 Limitations
Our method still inherits the problems of the baseline AnyText, such as the unsatisfactory ability to generate texts on a small scale or to generate texts in a specific font. Restricted by the latent size, our method may not acquire accurate texts when the text region is too small.
Figure 8: Examples of texts blend perfectly with background. Visual text generated by AnyText in the first row is often blurry and wrong. While our method can not only generate accurate visual text but the visual text is also blended with the background perfectly.
Figure 9: User study print screen.
Figure 10: Diversity demonstration.The first row demonstrates various fonts and colors a single prompt can generate. The second row shows same texts with different prompts can still generate coherent and reasonable images. We use different masks in the third row. The fourth row illustrates the ability to simply change the language of the generated texts.
Figure 11: Additional slanted results for our method.
Figure 12: Additional curved results for our method.
Figure 13: Additional complex results for our method.

### v2

@@ 4.3 Quantitative Analysis
As shown in Tab. 2, our method shows substantial improvements over the competitive methods in both English and Chinese across all levels. Notably, our method doesn’t require additional training and enhances the performance of Diff-Text, GlyphControl, and AnyText at the easy level. At the hard level, AnyText with our method yields results closest to the ground truth in the NED metric, with an improvement of approximately $10\%.$ over the baseline AnyText in both languages. As similarly shown in Tab. 3, our methods outperform baselines on vanilla AnyText Benchmark except FID. Meanwhile, Diff-Text with our method shows an impressive $20\%$ improvement at the hard level in the English set.
@@
In the absence of ground-truth images, we use the CLIP score [11] to assess the consistency between the prompt and the generated image. We compute the average cosine similarity between the prompt and the generated image, excluding the influence of the visual texts. As shown in Tab. 2, our method increases the accuracy of visual text generation without compromising the baseline performance. Additionally, our method slightly improves text-image similarity due to our design, which minimizes conflicts between the image composition and the visual text. This leads to a more coherent layout for both the image and the text.
@@
Following [2], we conduct a user study to comprehensively compare the generation results in Tab. 4. We create 27 input sets of varying difficulty. For each set, participants receive our input prompt, position mask, and two images (our result and a baseline) in random order, and select the image with the best quality and most accurate visual text. The final score is the average number of selections per prompt. We gather 168 judgments from a diverse group of experts and non-experts and report vote percentages. As shown in Tab. 4, our method is preferred in all cases.
Figure 6: More results on Background coherence and complex text layouts. The left two columns show how our approach seamlessly blends text with the background, while the remaining images highlight its ability to produce multi-sentence and circular text.
4.4 Qualitative Comparisons
As presented in Fig. 5, TextDiffuser [6] tends to generate low-quality images, specifically in the second and the fifth rows. Textdiffuser-2 [5] fails to consistently generate legible text within specified bounding boxes. We suspect this occurs due to interference between textual and non-textual elements when using overly large bounding boxes. GlyphControl [38] frequently produces texts outside its designated area, particularly when the text is near the image boundary, as shown in the second and fourth row. SceneVTG [43] suffers from severe text distortion, especially in the first and third rows. Background occlusion is also evident in the second, third, and fourth rows, where the texts “Burning,” “Mamba,” and “Twinkle” are obscured by lava, a human figure, and stars, respectively. This suggests SceneVTG struggles with complex backgrounds and lacks seamless integration with other text-to-image models. TextHarmony [42] faces similar issues. In image editing mode, it struggles to properly fill blanks with complex shapes, leading to distorted text, as seen in the second row. For Diff-text [41], there is a noticeable lack of coherence in the images due to insufficient consideration for harmony between the visual texts and the background. AnyText [36] tends to produce distorted visual texts due to a loss of semantic and structural information during inference. For example, in the first row, the text “Track” overlaps with the train head, compromising both elements. In contrast, our method generates high-quality images with accurately rendered English and Chinese texts—even in challenging scenarios such as those in the fourth and fifth rows and Fig. 6. Moreover, our approach not only corrects text distortion but also reduces background occlusion between textual and non-textual elements. For instance, in the third row of Fig. 5, AnyText [36] disrupts the text “Mamba” with a human head, whereas our method repositions the head to produce clear text alongside well-integrated non-text elements.
Figure 7: Ablation Study Visualization. The first column shows accuracy improvement with SRB. The second reveals that SIB enhances accuracy but compromises coherence. The third confirms that combining both branches enhances both. Finally, the fourth shows that adding AdaIN further improves image quality.
@@ 5 Ablation Study
In this section, we demonstrate the effectiveness of our components through an ablation study on the medium level. We use OCR accuracy (Sen.Acc and NED) and CLIP score as the main metrics, given their importance in evaluating visual text generation. All parameters are kept consistent with those outlined in Sec. 4.1.
@@
Table 5: Ablation study.
@@ 6 Conclusion
We advance visual text generation by tackling the challenge of complex text synthesis. Our proposed STGen introduces a dual-branch approach: the Semantic Rectification Branch, which refines text generation using latent extracted from simpler scenarios, and the Structure Injection Branch, which enhances text structure by incorporating latent of glyph image. For highly challenging cases, we break them into more manageable cases. Integrated via a dedicated ControlNet, STGen seamlessly enhances existing models. Extensive experiments on our benchmark confirm its superior performance, making STGen a promising step toward real-world applications.